In a large restaurant 2 out of 5 customers ask for water with their meals .A random sample of 12 customers is selected . Find the probability that exactly 8 customers ask for water with their meal.


1
Expert's answer
2021-12-21T15:15:32-0500

binomial distribution

"n=12,p=2\/5=0.4,q=1-p=0.6"

"P(x=k)=C^k_{n}\cdot p^k\cdot q^{-n-k}"


"P(x=8)=C^8_{12}\cdot0.4^8\cdot0.6^4=0.0420"